1.A card is drawn from a standard deck of 52 cards and then placed back into the deck. Find the probability that a red card is drawn at least once by the third draw.
Answer by greenestamps(13200)



The opposite of having at least one red card in three draws is having no red cards in three draws.

P(no red cards in three draws) = (1/2)(1/2)(1/2) = 1/8.

P(at least one red card in three draws) = 1-1/8 = 7/8.